A gasoline engine with a thermal efficiency of 28% has a power out put of 50hp . How much heat must be supplied per second to the engine
1
Expert's answer
2021-08-23T12:41:48-0400

By definition, the efficiency is given as follows:


"\eta = \dfrac{P_{out}}{P_{in}}\times 100\%"

where "P_{out} = 50hp" is the output power of the engine, and "P_{in }" is the heat supplied per second to the engine. Thus, obtain:


"P_{in} = \dfrac{P_{out}}{\eta}\times 100\% = \dfrac{50hp}{28\%}\times 100\% \approx 179hp"

Answer. 179 hp.